Preliminary figures (in billions of dollars) for 2004 taken from the 2005 Economic Report of the President showed that: Y-1 1,728.0, С 8,231.1. EX-I, 170.2, IM 1.779.6, G-2,183.8 and T 1771.7. We can use the national income identities to describe the financing of U.S. business investment and the U.S. government deficit Part A: Use the income identities to find what U.S. private business investment, I, was in 2004 Show your work. Part B: Use the income identities to find what U.S. private savings, S, was in 2004. Show your work. Part C: Both U.S. private business investment and the U.S. government deficit need to be financed each year. What were the total financing needs of business and government in 2004? Part D: Was private savings sufficient to finance both private business investment and the government deficit? What was the shortfall? Where did U.S. firms and the U.S. government find the extra financing? Can you see the shortfall in the numbers provided at the outset (Hint: think about the current account)?
Answer #1

1) Solution: U.S. private business investment in 2004: 1,922.40

Working:

Y = C + I + G + EX - IM

I = [Y - (C + G + EX - IM)]

I = [11,728.0 - (8,231.1 + 2,183.8 + 1,170.2 - 1,779.6)]

I = 1,922.40

2) Solution: U.S. private savings in 2004: 1725.2

Working:

Y = C + S + T

S = [Y - (C + T)]

S = [11,728.0 - (8,231.1 + 1771.7)]

S = 1725.20

3) Solution: Total financing requirements of business and government in 2004: 2,334.50

Working:

I + (G - T) = [1,922.4 + (2,183.8 - 1771.7)]

= 2334.50

4) Solution: The private savings were insufficient to finance both private business investment as well the government deficit due to the shortfall of 609.30. Consequentky had to borrow from foreigners. The current account deficit depicts the U.S. borrowing from abroad.

Working:

Shortfall = [I + (G - T)] - S

Shortfall = [2334.5 - 1725.2]

Shortfall = 609.30

or;

[I + (G - T)] - S = - current account = [IM - EX]

= [1,779.60 - 1,170.20]

= 609.3